Find the standard normal area for each of the following (Round your answers to 4 decimal places.):
Standard normal area
a. P(1.22 < Z < 2.12)      
b. P(2.01 < Z < 3.01)      
c. P(-2.01 < Z < 2.01)      
d. P(Z > 0.51)      
Answer #1

a)

P( 1.22 < Z < 2.12) = P( Z < 2.12 ) - P( Z < 1.22)

= 0.983 - 0.8888

= 0.0942

b)

P( 2.01 < Z < 3.01) = P( Z < 3.01) - P( Z < 2.01)

= 0.9987 - 0.9778

= 0.0209

c)

P( -2.01 < Z < 2.01) = P( Z < 2.01) - p( Z < -2.01)

= 0.9778 - ( 1 - 0.9778)

= 0.9556

d)

P( Z > 0.51) = 1 - P( Z < 0.51)

= 1 - 0.6950

= 0.3050